# Rotating twisted magnetosphere of magnetars: approximate analytical solutions

### v1

@@ 2.3 The particle outflow luminosity
$L=2\int_{0}^{\Psi_{0}}Id\Psi\propto\Psi_{0}^{2},$ (17)
where $I$ is the outflowing current, $\Psi_{0}$ the open field flux. From the definition of toroidal field, the relation between $I$ and $\Phi$ is: $I=(c/2)\Phi$. Using the solution for $\Phi$ (eq.4), the particle outflow luminosity is proportional to $\Psi_{0}^{2}$. It can be shown that the above definition is equal to the Poynting flux in the open field line regions (Stefanou et al. 2023). Another point of view is: this definition is actually the work done on the outflowing particles. The current is proportional to $\Phi$, the maximum acceleration potential is proportional to $\Psi$. Therefore, the power defined in this way is just the current times the voltage. Therefore, we’d like to call it the “particle outflow luminosity". Of course, here the “particle outflow" means both the outflowing particles and the electromagnetic fields associated with them.
There is a tricky point here. In the case of normal pulsars, the rotational energy is the sole energy source. Therefore, the particle outflow luminosity can be named as the rotational energy loss rate. In the case of magnetars, the magnetic energy release may dominate over the rotational energy loss rate, e.g. magnetar’s persistent X-ray luminosity can be higher than their rotational energy loss rate. In this case, the particle outflow luminosity may also be powered by the magnetic energy release. Noting this point, we just call it particle outflow luminosity.
@@
From the above express, it can be seen that the particle outflow luminosity for a rotating twisted magnetosphere can be much larger than the rotating dipole case. This may explain the enhanced spin-down during magnetar outburst (Archibald et al. 2020; Huang et al. 2023). Furthermore, during the outburst, the magnetosphere may evolve towards the dipole case, i.e. $n$ evolves from some initial value $n_{0}$ towards $n=1$. Then the spin-down torque will decrease with time during the outburst. This is consistent with the general trend of spin-down evolution during outburst.
Here the twist of the field line is characterized by the parameter $n$. Geometrically, the twist of the field line along the axis from the north pole to the south pole is defined as the maximum twist $\Delta\phi_{\rm max}$ (Thompson et al. 2002). There is a one-to-one relation between the parameter $n$ and the maximum twist (Pavan et al. 2009; Tong 2019):
@@
During the numerical simulations (Ntotsikas et al. 2024, 2025), the twist defined there is half the value reported in previous works. Therefore, in order to compare with the numerical simulations, half the maximum twist should be used:
@@
Then the Y-point radius, polar cap, and particle outflow luminosity can all be obtained as a function of twist $\Delta\phi$. This is also what is reported in the numerical simulations.
@@ 3 Comparison with the numerical simulations
Previous numerical simulations have not include the twist of the magnetic field (Contopoulos et al. 1999; Spitkovsky 2006). In Ntotsikas et al. (2024, 2025), the twist in the closed field line regions are included. While the twist in the open field line regions is still due to the rotation of the neutron star. In the numerical simulations, it is found that (Ntotsikas et al. 2024, 2025): (1) The Y-point radius is smaller than the light cylinder radius, (2) The polar cap is larger compared with the dipole case, (3) the particle outflow luminosity is higher. Point (2) & (3) are already known in the analytical result in Tong (2019). Here, all three points are obtained analytically (see the above section). The analytical solutions will be compared with the numerical simulations.
In the numerical simulations, the toroidal field (or poloidal current) in the closed field lines are prescribed in a power-law form. While, in our approximate analytical solutions, the closed field line region is assumed to be in a self-similar form. Despite this difference, the maximum twist can be defined in both cases. Therefore, we will compare the Y-point radius, polar cap, and particle outflow luminosity as a function of maximum twist. Of course, the twist reported in the numerical simulations is actually half the maximum value.
In table 1-3 in Ntotsikas et al. (2024), table 1 in Ntotsikas & Gourgouliatos (2025), the numerical results of open field line flux $\Psi_{0}$, normalized particle outflow luminosity compared with the dipole case $L_{\rm twisted}/L_{\rm dipole}$, Y-point radius $R_{Y}$ and half the maximum twist $\Delta\phi$ are shown. In the presence of twist, the field line will be different from the dipole case, the polar cap should be found by tracing back the field line to the neutron star surface (Ntotsikas & Gourgouliatos 2025). As an approximation, the flux function will be like the dipole case at the neutron star surface (eq.(19) & (20) in Ntotsikas et al. 2024. Actually, it is one of the boundary conditions in the numerical simulations. This expression holds for both the dipole and twisted dipole case, as has been proved above.):
@@
The comparison between our approximate analytical solutions and the numerical simulations are shown in figure 1, 2, and 3. From this comparison, it can be seen that: (1) the general trend of both the analytical approximations and the numerical results are consistent with each other. (2) For the particle outflow luminosity, the analytical results and the numerical simulations of Ntotsikas & Gourgouliatos (2025) are consistent with each other (magenta triangle in figure 3). We consider these as two pieces of evidence in support of our approximate analytical solutions.
@@ 4 Discussions
From the magnetohydrodynamics point of view, the magnetic field structure (eq.5 and 4) forms the basics of all subsequent calculations. Here we want to focus on the radio emission of magnetars. The fast radio burst may also be related with the magnetospheric physics of magnetars.
The polar cap (eq.16) is related directly to the observed hot spot during magnetar outburst. During the outburst, an untwisting magnetosphere will result in a shrinking hot spot (for the parameter $n$ evolves from $n_{0}$ to 1). At the same time, the maximum acceleration potential across the polar cap is closed related to the turn-on/off of radio emission (Ruderman & Sutherland 1975):
@@
For a given parameter $0<n<1$, the corresponding maximum acceleration potential can be obtained. In order to obtain the evolution of magnetosphere, the evolution of the twist should be calculated (i.e., the evolution of $n$). This will depend on the particle acceleration mechanism etc (Beloborodov 2009; Tong 2019). At the same time, the magnetar X-ray flux may decay with time in an exponential form in observations (Coti Zelati et al. 2018). Therefore, the magnetic free energy stored in the twisted magnetic field may also decay with time in an exponential form. Assuming this exponential form of evolution, a simplified solution can be found for the parameter $n$ (Tong & Huang 2020):
@@
Using this approximation, a typical evolution of the maximum acceleration potential of a magnetar can be obtained, in figure 4. During the calculations, a typical period of $5\ \rm s$ and polar magnetic field strength $B_{p}=10^{12}\ \rm G$ is assumed. The initial value and decaying timescale of $n$ is: $n_{0}=0.5$, $\tau=1\ \rm yr$. Figure 4 demonstrate that if a magnetar’s true dipole field is only about $10^{12}\ \rm G$ (its spin-down rate may be greatly enhanced by the particle outflow luminosity, therefore its true magnetic field may be much smaller than the characteristic magnetic field), during the outburst its polar cap and maximum acceleration potential may be greatly enhanced. The maximum acceleration potential may exceed $10^{12}\ \rm V$, the threshold value for typical radio emission. During the outburst decaying phase, its acceleration potential decays with time. Later, it may even fall below the radio emission threshold. Then the radio emission may turn off. This may be one reason for the radio turn-on/off of magnetars.
@@
This definition of pulse width is in the range 010-1. Therefore, it is actually the duty cycle of the pulse profile. Then for a typical magnetar with period $P=5\ \rm s$, the pulse width as a function of emission height for several values of $n$ are given in figure 5. For a larger twist (smaller $n$), the pulse width will be larger. Compared with normal pulsars, the radio emission of magnetar typically have wide pulse profile, by noting that magnetar have a long pulsation period (Camilo et al. 2006, 2008; Levin et al. 2012; Huang et al. 2021). From figure 5, for a typical pulse width about $10\%$ of magnetars, (1) a very high emission height is required for a dipole field ($n=1$), (2) for a twisted magnetosphere, the pulse width can be explained by an emission about 100 neutron radius (the typical value of normal pulsars, Johnston & Kramer 2019). During the outburst decaying phase, the pulse width decreases with time. It is possible that at a later time, the line of sight will be out-of the radio emission beam. This may be another reason for the radio trun-on/off of magnetar radio emission.
@@ 5 Conclusions
Comparisons with the numerical simulations are made. The general trend of the approximate analytical solution and the numerical simulations are the same. This may be viewed as support for the analytical approximation. The analytical solutions can be used to model the evolution of the magnetar magnetosphere. We have only showed the case of evolving maximum acceleration potential, and pulse width. These two points may be related to the radio turn-on/off of magnetars. In our opinion, the modeling of radio emission in the magnetar magnetosphere (magnetars and fast radio burst) should be based on the rotating twisted magnetosphere, instead of the normal dipole case. In general, a rotating twisted magnetosphere has larger open field line regions. The radio emission of magnetars and fast radio bursts may both originate in the larger and evolving open field line regions of magnetars.
@@ Acknowledgements
This work is supported by National SKA Program of China (No. 2020SKA0120300, 2022SKA0120102) and NSFC (12133004, 12173066).

### v2

@@ 2.3 The particle outflow luminosity
$L=2\int_{0}^{\Psi_{0}}\Phi d\Psi\propto\Psi_{0}^{2},$ (17)
where $\Psi_{0}$ is the flux associated with the last open field line. From its definition, $\Phi$ denotes the poloidal current (up to a numerical factor). Using the solution for $\Phi$ (eq.4), the particle outflow luminosity is proportional to $\Psi_{0}^{2}$. It can be shown that the above definition is equal to the Poynting flux in the open field line regions (Stefanou et al. 2023). Another point of view is: this definition is actually the work done on the outflowing particles. The current is proportional to $\Phi$, the maximum acceleration potential is proportional to $\Psi$. Therefore, the power defined in this way is just the current times the voltage. Therefore, we’d like to call it the “particle outflow luminosity". Of course, here the “particle outflow" means both the outflowing particles and the electromagnetic fields associated with them.
There is a tricky point here. In the case of normal pulsars, the rotational energy is the only energy source. Therefore, the particle outflow luminosity can be named as the rotational energy loss rate. In the case of magnetars, the magnetic energy release may dominate over the rotational energy loss rate, e.g. magnetar’s persistent X-ray luminosity can be higher than their rotational energy loss rate. In this case, the particle outflow luminosity may also be powered by the magnetic energy release. Noting this point, we just call it particle outflow luminosity.
@@
From the above express, it can be seen that the particle outflow luminosity for a rotating twisted magnetosphere can be much larger than the rotating dipole case. This may explain the enhanced spin-down during magnetar outburst (Archibald et al. 2020; Huang et al. 2023). Furthermore, during the outburst, the magnetosphere may evolve towards the dipole case, i.e. $n$ evolves from some initial value $n_{0}$ towards $n=1$. Then the spin-down torque will decrease with time during the outburst. This is consistent with the general trend of spin-down evolution during magnetar outburst.
Here the twist of the field line is characterized by the parameter $n$. Geometrically, the twist of the field line from the north pole to the south pole is defined as the maximum twist $\Delta\phi_{\rm max}$ (Thompson et al. 2002). There is a one-to-one relation between the parameter $n$ and the maximum twist (Pavan et al. 2009; Tong 2019):
@@
The maximum allowed twist is $\pi$ radians for a twisted dipole field (Lynden-Bell & Boily 1994; Wolfson 1995; Thompson et al. 2002). The analytical result of the self-similar solution (eq.19) can catch this point approximately. And the analytical result agrees with the detailed calculations, especially for small twist case (Pavan et al. 2009; Tong 2019). During the numerical simulations (Ntotsikas et al. 2024, 2025), the twist defined there is half the value reported in previous works. Therefore, in order to compare with the numerical simulations, half the maximum twist should be used:
@@
Then the Y-point radius, polar cap, and particle outflow luminosity can all be obtained as a function of twist $\Delta\phi$.
@@ 3 Comparison with the numerical simulations
Previous numerical simulations have not include the twist of the magnetic field (Contopoulos et al. 1999; Spitkovsky 2006). Gruzinov (2006) had discussed that adding twist to the field lines may result in the opening of more field lines. In Ntotsikas et al. (2024, 2025), the twist in the closed field line regions is included, while the twist in the open field line regions is still due to the rotation of the neutron star. In the numerical simulations, it is found that (Ntotsikas et al. 2024, 2025): (1) The Y-point radius is smaller than the light cylinder radius, (2) The polar cap is larger compared with the dipole case, (3) the particle outflow luminosity is higher. Point (2) & (3) are already known in the analytical result in Tong (2019). Here, all three points are obtained analytically (see the above section). The analytical solutions will be compared with the numerical simulations.
In the numerical simulations, the toroidal field (or poloidal current) in the closed field lines are prescribed in a power-law form. While, in our approximate analytical solutions, the closed field line region is assumed to be in a self-similar form. Despite this difference, the maximum twist can be defined in both cases. Therefore, we will compare the Y-point radius, polar cap, and particle outflow luminosity as a function of the maximum twist. Of course, the twist reported in the numerical simulations is actually half the maximum value.
In table 1-3 in Ntotsikas et al. (2024), table 1 in Ntotsikas & Gourgouliatos (2025), the numerical results of the flux associated with the last open field line $\Psi_{0}$, normalized particle outflow luminosity compared with the dipole case $L_{\rm twisted}/L_{\rm dipole}$, Y-point radius $R_{Y}$ and half the maximum twist $\Delta\phi$ are shown. In the presence of twist, the field line will be different from the dipole case, the polar cap should be found by tracing back the field line to the neutron star surface (Ntotsikas & Gourgouliatos 2025). As an approximation, the flux function will be like the dipole case at the neutron star surface (eq.(19) & (20) in Ntotsikas et al. 2024. Actually, it is one of the boundary conditions in the numerical simulations. This expression holds for both the dipole and twisted dipole case, as has been proved above.):
@@
The comparisons between our approximate analytical solutions and the numerical simulations are shown in figure 1, 2, and 3. From this comparison, it can be seen that: (1) the general trend of both the analytical approximations and the numerical results are consistent with each other. (2) For the particle outflow luminosity, the analytical results and the numerical simulations of Ntotsikas & Gourgouliatos (2025) are consistent with each other (magenta triangle in figure 3). We consider these as two pieces of evidence in support of our approximate analytical solutions.
@@ 4 Discussions
From the magnetohydrodynamics point of view, the magnetic field structure (eq.5 and 4) forms the basis of all subsequent calculations. Equation (5) is the result of self-similar twisted dipole field. In the closed field line region, the relation between the current and flux is non-linear (Wolfson 1995). While, eq.(4) represents a linear relation between the current and flux in the open field line regions, and it agrees with the minimum torque model. In general, the relation between the current and flux can be non-linear. The poloidal field is unchanged. Only the toroidal field changes for a different current. (1) For a split monopole solution (Contopoulos 2005): $\Phi=-\Omega\Psi(2-\Psi/\Psi_{\rm open})$, it is consistent with eq.(4) for small $\theta$ (near the magnetic pole). For the toroidal field at the Y-point ($\Psi=\Psi_{\rm open}$), it is different from eq.(4) by a factor of two. Since there is an undetermined numerical factor during the definition of the Y-point radius, the final express for the Y-point radius will be the same (eq.(12)). (2)For other non-linear forms between the current and flux, the toroidal field, Y-point radius etc may be calculated similarly. However,
  1. 1.
    we are not sure whether a non-linear relation between the current and flux is still an approximate solution to the original pulsar equation.
  2. 2.
    Irrespective of the details, from previous experiences of analytical and numerical calculations (Tong 2019; Ntotsikas et al. 2024), the general trend should be the same: more twist will result in more open field lines, and consequently larger polar cap and particle outflow luminosity.
Here we want to discuss the radio emission of magnetars. The fast radio bursts may also be related with the magnetospheric physics of magnetars. The polar cap (eq.16) is related directly to the observed hot spot during magnetar outburst. During the outburst, an untwisting magnetosphere will result in a shrinking hot spot (for the parameter $n$ evolves from $n_{0}$ to 1). At the same time, the maximum acceleration potential across the polar cap is closed related to the turn-on/off of radio emission (Ruderman & Sutherland 1975):
@@
For a given parameter $0<n<1$, the corresponding maximum acceleration potential can be obtained. In order to obtain the evolution of magnetosphere, the evolution of the twist should be calculated (i.e., the evolution of $n$). This will depend on the particle acceleration mechanism (Beloborodov 2009; Tong 2019). At the same time, the magnetar X-ray flux may decay with time in an exponential form in observations (Coti Zelati et al. 2018). Therefore, the magnetic free energy stored in the twisted magnetic field may also decay with time in an exponential form. Assuming this exponential form of evolution, a simplified solution can be found for the parameter $n$ (Tong & Huang 2020):
@@
Using this approximation, a typical evolution of the maximum acceleration potential of a magnetar can be obtained, in figure 4. During the calculations, a typical period of $5\ \rm s$ and polar magnetic field strength $B_{p}=10^{12}\ \rm G$ is assumed. The initial value and decaying timescale of $n$ is: $n_{0}=0.5$, $\tau=1\ \rm yr$. Figure 4 demonstrate that if a magnetar’s true dipole field is only about $10^{12}\ \rm G$ (its spin-down rate may be greatly enhanced by the particle outflow luminosity, therefore its true magnetic field may be much smaller than the characteristic magnetic field), during the outburst its polar cap and maximum acceleration potential may be greatly enhanced. The maximum acceleration potential may exceed $10^{12}\ \rm V$, the threshold value for typical radio emission. During the outburst decaying phase, its acceleration potential decays with time. Later, it may even fall below the radio emission threshold. Then the radio emission may turn off. This may be the first reason for the radio turn-on/off of magnetars.
@@
This definition of pulse width is in the range 010-1. Therefore, it is actually the duty cycle of the pulse profile. Then for a typical magnetar with period $P=5\ \rm s$, the pulse width as a function of emission height for several values of $n$ are given in figure 5. For a larger twist (smaller $n$), the pulse width will be larger. Compared with normal pulsars, the radio emission of magnetar typically have wide pulse profile, by noting that magnetar have a long pulsation period (Camilo et al. 2006, 2008; Levin et al. 2012; Huang et al. 2021). From figure 5, for a typical pulse width about $10\%$ of magnetars, (1) a very high emission height is required for a dipole field ($n=1$), (2) for a twisted magnetosphere, the pulse width can be explained by an emission height about 100 neutron star radius (the typical value of normal pulsars, Johnston & Kramer 2019). During the outburst decaying phase, the pulse width decreases with time. It is possible that at a later time, the line of sight will be out-of the radio emission beam. This may be the second reason for the radio turn-on/off of magnetars.
@@ 5 Conclusions
Comparisons with the numerical simulations are made. The general trend of the approximate analytical solution and the numerical simulations are the same. This may be viewed as support for the analytical approximation. The analytical solutions can be used to model the evolution of the magnetar magnetosphere. We have only showed the case of evolving maximum acceleration potential, and pulse width. These two points may be related to the radio turn-on/off of magnetars. In our opinion, the modeling of radio emission in the magnetar magnetosphere (magnetars and fast radio bursts) should be based on the rotating twisted magnetosphere, instead of the normal dipole case. In general, a rotating twisted magnetosphere has larger open field line regions. The radio emission of magnetars and fast radio bursts may both originate in the larger and evolving open field line regions of magnetars.
@@ Acknowledgements
This work is supported by NSFC (12133004, 12173066) and National SKA Program of China (No. 2020SKA0120300, 2022SKA0120102).